### v1

@@ 3.2 Compact semi-simple algebras
For a full exposition with proofs of the facts quoted in the previous remark, see Appendix B in [27] or Knapp’s book [18, Theorem 6.11], where we picked $h_{\alpha}=iH_{\alpha}$, and
@@
We remark that for $U(n)$ we established the existence of norming functionals adapted to a vector $v$ by permutating the elements of the basis of $\mathbb{C}^{n}$ [26, Lemma 2.38]. With the same idea, the proof can be extended to $SU(n)$ and $O(n)$. However, since in general the group $G$ might not act transitively on the positive roots (roots might even have different lengths), this mechanism of proof is not suitable for generalization. In what follows we establish in general the existence of norming functionals adapted to $v$, but with an entirely different mechanism of proof.
@@ 3.3 The form $S$ and sectional curvature
If $G=SU(2)$, and the eigenvalues of yxy-x are equal then $y-x=\lambda 1$, hence [x,y]=[x,yx]=0[x,y]=[x,y-x]=0 and all the conditions of the previous theorem are equivalent. Otherwise yxy-x is regular and again all the conditions of the previous theorem are equivalent. This indicates a miscalculation in [3, Example 4.4]; that $\rho$ has $\|\rho\|_{1}=2.8\neq 1$.
@@ 3.3.1 Sectional curvature
where $\dim(H)=\dim(\operatorname{\mathfrak{g}})+1=d+1$. Using the Cartan-Alexandrov-Toponogov Comparison Theorem [11, Theorem 6.5.6], by fixing the lengths of two sides of the triangle in $G$ and the angle substended between them, it follows that

### v2

@@ 3.2 Compact semi-simple algebras
For a full exposition with proofs of the facts quoted in the previous remark, see Appendix B in [22] or Knapp’s book [15, Theorem 6.11], where we picked $h_{\alpha}=iH_{\alpha}$, and
@@
We remark that for $U(n)$ we established the existence of norming functionals adapted to a vector $v$ by permutating the elements of the basis of $\mathbb{C}^{n}$ [21, Lemma 2.38]. With the same idea, the proof can be extended to $SU(n)$ and $O(n)$. However, since in general the group $G$ might not act transitively on the positive roots (roots might even have different lengths), this mechanism of proof is not suitable for generalization. In what follows we establish in general the existence of norming functionals adapted to $v$, but with an entirely different mechanism of proof.
@@ 3.3 The form $S$ and sectional curvature
If $G=SU(2)$, and the eigenvalues of yxy-x are equal then $y-x=\lambda 1$, hence [x,y]=[x,yx]=0[x,y]=[x,y-x]=0 and all the conditions of the previous theorem are equivalent. Otherwise yxy-x is regular and again all the conditions of the previous theorem are equivalent. This indicates a miscalculation in [2, Example 4.4]; that $\rho$ has $\|\rho\|_{1}=2.8\neq 1$.
@@ 3.3.1 Sectional curvature
where $\dim(H)=\dim(\operatorname{\mathfrak{g}})+1=d+1$. Using the Cartan-Alexandrov-Toponogov Comparison Theorem [9, Theorem 6.5.6], by fixing the lengths of two sides of the triangle in $G$ and the angle substended between them, it follows that